# Dept. of Systems Engineering, ETS Montreal, Canada

### v1

@@ 0.E.3 Part-prototype masking:
Figure 0.E.4: Prototypes regions extracted by the PRM module for (a) infrared and (b) visible images. Note that the mask size is 18$\times$9, which is then resized to fit the original input image. As shown, the mask of prototypes focuses on similar body parts without accounting for identity.
0.E.4 Semi-supervised body part detection:
An additional benefit of our HCL module lies in its ability to detect meaningful parts in a semi-supervised manner. By forcing the model to identify semantic regions that are both informative about foreground objects and contrastive to each other, our hierarchical contrastive learning provides robust part detection, even in the absence of part labels. To assess HCL, we fine-tuned our ReID model as a student using a pre-trained part detector [part_detec] on the PASCAL-Part Dataset [chen2014detect] as the teacher. In Fig. 0.E.5, the results show our model’s strong capacity for detecting body parts compared to its teacher.
@@ 0.E.4 Semi-supervised body part detection:
Figure 0.E.5: Semi-supervised part discovery on the SYSU-MM01 dataset. The first columns are the (a) infrared and (b) visible images. The second column images are the result from [part_detec], and the last column are results with our fine-tuned model in BMDG.

### v2

@@ E.3 Part-prototype masking:
Figure E.4: Prototypes regions extracted by the PRM module for (a) infrared and (b) visible images. Note that the mask size is 18$\times$9, which is then resized to fit the original input image. As shown, the mask of prototypes focuses on similar body parts without accounting for identity.
E.4 Semi-supervised body part detection:
An additional benefit of our HCL module lies in its ability to detect meaningful parts in a semi-supervised manner. By forcing the model to identify semantic regions that are both informative about foreground objects and contrastive to each other, our hierarchical contrastive learning provides robust part detection, even in the absence of part labels. To assess HCL, we fine-tuned our ReID model as a student using a pre-trained part detector [part_detec] on the PASCAL-Part Dataset [chen2014detect] as the teacher. In Fig. E.5, the results show our model’s strong capacity for detecting body parts compared to its teacher.
@@ E.4 Semi-supervised body part detection:
Figure E.5: Semi-supervised part discovery on the SYSU-MM01 dataset. The first columns are the (a) infrared and (b) visible images. The second column images are the result from [part_detec], and the last column are results with our fine-tuned model in BMDG.